OpenStudy (anonymous):

solve for the unknown 216=36^3x

OpenStudy (stormfire1):

Is that \[36^3x\] or \[36^{3x}\]

OpenStudy (anonymous):

second choice

OpenStudy (anonymous):

\[6^3=(6^2)^{3x}\] is a start

OpenStudy (stormfire1):

Danyel, do you understand it from there?

OpenStudy (anonymous):

I guess not because when I put that in I end up with a decimal.

OpenStudy (anonymous):

My book doesnt even tell me how to do this.

OpenStudy (stormfire1):

On the left you have 6^3 and you need to make the right equal that. On the right, you have \[(6^2)^{3x}\]. Using the rules of exponents, you should see that this is an exponent raised to an exponent. That means you need to multiply the exponents together and ultimately they should = the "3" that's on the left side of the equation. 2 * 3x = 3 x must be 1/2

OpenStudy (anonymous):

why won't it be 6 because 2*3=6

OpenStudy (stormfire1):

x = 1/2 so: 2 * 3 * 1/2 = 3

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

glad I asked

OpenStudy (stormfire1):

You can of course check your answer using a calculator and plugging in 1/2 for x in the original equation. You'll see that it = 216

OpenStudy (anonymous):

so when I write out the answer I will write x=1/2?

OpenStudy (stormfire1):

Yes

OpenStudy (anonymous):

I think I am starting to understand a bit better. thank you

OpenStudy (stormfire1):

glad to help